OpenStudy (anonymous):

Solve: 7(x − 1) = 5(x + 2) + 1

OpenStudy (anonymous):

7 (x-1) = 5 (x+2)+1 Expand out terms on both sides: 7 x-7 = 5 x+11 Add (7-5 x) to both sides: 2 x = 18 Divide both sides by 2: x = 9